# Dynamic Attention Analysis for Backdoor Detection in Text-to-Image Diffusion Models

### v1

@@ III-B Anomalous Cue of Backdoor Samples
The result implies a phenomena that the attention of the $<$EOS$>$ token in backdoor samples dissipate slower than the one in benign samples. Recall that the $<$EOS$>$ serves as the summary token that carries most information of the prompt [29]. Intuitively, for a benign sample, attention gradually shifts from focusing on the $<$EOS$>$ token to important tokens throughout the denoising process. This results in the $<$EOS$>$ token having a higher rate of attention evolve compared to the average of other tokens. In contrast, for a backdoor sample, since the model is forced to learn a specific pattern, attention spreads more slowly to the context. As a result, the $<$EOS$>$ token exhibits a lower rate of attention evolve compared to the average of other tokens.
@@ V Conclusion
Theorem 1. If the square matrix $B=[b_{ij}]_{i,j=1}^{n}$ is a real matrix and satisfies $b_{ii}>\sum_{j\neq i}^{n}|b_{ij}|,\quad(i=1,2,\dots,n),$ then $B$ is non-singular thus $\text{det}B>0$ [39].
Theorem 2. If the square matrix $B=[b_{ij}]_{i,j=1}^{n}$ is a real symmetric matrix and all of its eigenvalues are real, then there exists an orthogonal matrix $Q$ such that $Q^{T}BQ=Q^{-1}BQ$ is a diagonal matrix [39].
@@
where $X(t)\in\mathbb{R}^{n}$ is the state vector, and $f(\cdot):\mathbb{R}^{n}\times T\to\mathbb{R}^{n}$ is a function of the state and time. If there exists a positive definite, continuously differentiable scalar function $V[X(t),t]$ in the neighborhood of the equilibrium state $X^{*}$, with its parameter $V[X(t),t]$ being negative definite, and if as
@@
For the first term, since $F$ is a symmetric diagonal matrix and its diagonal elements $\gamma_{i}$ satisfy $\gamma_{i}<0$, it follows that for $X(t)\neq X_{e}$:
@@
where,
$\gamma_{\max}=\max\{\gamma_{1},\gamma_{2},\dots,\gamma_{n}\}<0,$ (30)
thus, $X^{T}(t)(F^{T}+F)X(t)<0$.
For the second term, let matrix $A^{\prime}=[a^{\prime}_{ij}]_{i,j=1}^{n}$ and define $A^{\prime}=A^{T}+A$. Since $A^{\prime}$ is a symmetric matrix, all its eigenvalues are non-positive real numbers. This follows from:
$\sum_{j=1}^{n}a_{ij}=0\quad(i=1,2,\dots,n),$ (31)
and
$\sum_{i=1}^{n}a_{ij}=0\quad(j=1,2,\dots,n).$ (32)
There exists an $n$-dimensional vector $Y^{\prime}=(1/\sqrt{n},1/\sqrt{n},\dots,1/\sqrt{n})$ satisfying:
$A^{\prime}Y^{\prime}=0Y^{\prime},$ (33)
which implies that 0 is an eigenvalue of matrix $A^{\prime}$. Suppose $A^{\prime}$ has an eigenvalue $\lambda>0$, then:
$|\lambda I-A^{\prime}|=0,$ (34)
where $I$ is the identity matrix. In matrix $A^{\prime}$, any diagonal element satisfies:
$a^{\prime}_{ii}\leq 0,$ (35)
and any off-diagonal element satisfies:
$a^{\prime}_{ij}\geq 0\quad(i\neq j).$ (36)
Furthermore,
$a^{\prime}_{ii}=-\sum_{j=1,j\neq i}^{m}a^{\prime}_{i,j}.$ (37)
Define matrix $D=[d_{ij}]_{i,j=1}^{n}$, where $D=\lambda I-A^{\prime}$. Then, for any diagonal element of $D$, we have:
$d_{ii}=\lambda+|a^{\prime}_{ii}|.$ (38)
It is evident that:
@@
According to Theorem 1, we obtain:
$\text{det}D=\text{det}(\lambda I-A^{\prime})>0,$ (40)
so matrix $A^{\prime}$ does not have any positive real eigenvalues. Since $A^{\prime}$ is a symmetric matrix, its eigenvalues can be orthogonally diagonalized according to Theorem 2. Assuming the eigenvalues of $A^{\prime}$ are ordered as:
@@
and the corresponding orthogonal transformation matrix is $\Phi$, then:
$\begin{split}cX^{T}(t)(A^{T}+A)X(t)&=cX^{T}(t)A^{\prime}X(t)\\ &=c[\Phi X(t)]^{T}\operatorname{diag}\{\lambda_{1},\lambda_{2},\dots,\lambda_{n}\}\Phi X(t)\\ &=c\sum_{i=1}^{n}\lambda_{i}z_{i}^{2}(t),\end{split}$ (42)
where $\Phi X(t)=(z_{1}(t),z_{2}(t),\dots,z_{n}(t))^{T}$. Thus, when $X(t)\neq X_{e}$, we have:

### v2

@@ III-B Anomalous Cue of Backdoor Samples
The result implies a phenomena that the attention of the $<$EOS$>$ token in backdoor samples dissipate slower than the one in benign samples. Recall that the $<$EOS$>$ serves as the summary token that carries most information of the prompt [29, 63]. Intuitively, for a benign sample, attention gradually shifts from focusing on the $<$EOS$>$ token to important tokens throughout the denoising process. This results in the $<$EOS$>$ token having a higher rate of attention evolve compared to the average of other tokens. In contrast, for a backdoor sample, since the model is forced to learn a specific pattern, attention spreads more slowly to the context. As a result, the $<$EOS$>$ token exhibits a lower rate of attention evolve compared to the average of other tokens.
@@ V Conclusion
Theorem 1. Let $B=[b_{ij}]_{i,j=1}^{n}$ be a real square matrix. If it satisfies $b_{ii}>\sum_{j\neq i}^{n}|b_{ij}|,\quad(i=1,2,\dots,n),$ then $B$ is non-singular and in particular $\text{det}B>0$ [39].
Theorem 2. Let $B=[b_{ij}]_{i,j=1}^{n}$ be a real symmetric matrix. Then all its eigenvalues are real and there exists an orthogonal matrix $Q$ such that $Q^{T}BQ=Q^{-1}BQ$ is a diagonal matrix [39].
@@
where $X(t)\in\mathbb{R}^{n}$ is the state vector, and $f(\cdot):\mathbb{R}^{n}\times T\to\mathbb{R}^{n}$ is a function of the state and time. If there exists a continuously differentiable, positive definite scalar function $V[X(t),t]$ in the neighborhood of the equilibrium state $X^{*}$, with its parameter $V[X(t),t]$ being negative definite, and if as
@@
For the first term, since $F$ is a symmetric diagonal matrix with diagonal entries $\gamma_{i}$ satisfy $\gamma_{i}<0$, we have:
@@
[…]
For the second term, let $A^{\prime}=A^{T}+A$ and $A^{\prime}$ is a symmetric matrix. Under the assumptions:
$\sum_{j=1}^{n}a_{ij}=0\ \text{and}\ \sum_{i=1}^{n}a_{ij}=0,$ (30)
$A^{\prime}$ admits the eigenvalue $\lambda_{1}=0$ with eigenvector $Y^{\prime}=(1/\sqrt{n},\dots,1/\sqrt{n})^{T}$. If we suppose there exists a positive eigenvalue $\lambda>0$, then the matrix $D=\lambda I-A^{\prime}$ would satisfy:
$d_{ii}=\lambda+|a^{\prime}_{ii}|,\ d_{ij}=-a^{\prime}_{ij}\ (i\neq j),$ (31)
which leads to
@@
By Theorem 1, this implies $\text{det}D>0$, contradicting the assumption that $\lambda$ is an eigenvalue. Thus, all eigenvalues of $A^{\prime}$ are non-positive.
As $A^{\prime}$ is symmetric, it is orthogonally diagonalizable by Theorem 2. Let its eigenvalues of $A^{\prime}$ be ordered as:
@@
with corresponding orthogonal matrix $\Phi$, then:
$\begin{split}cX^{T}(t)A^{\prime}X(t)&=c[\Phi X(t)]^{T}\operatorname{diag}\{\lambda_{1},\lambda_{2},\dots,\lambda_{n}\}\Phi X(t)\\ &=c\sum_{i=1}^{n}\lambda_{i}z_{i}^{2}(t),\end{split}$ (34)
where $\Phi X(t)=(z_{1}(t),z_{2}(t),\dots,z_{n}(t))^{T}$. Since $\lambda_{i}\leq 0$, this term is also non-positive for $X(t)\neq X_{e}$. Combining both term yields: